The 2197-N block shown is in contact with 45° incline. The coefficient of static friction is 0.3. Compute the value of the horizontal force P necessary to just prevent motion down the incline.


1
Expert's answer
2021-03-02T18:09:41-0500

Let's apply the Newton's Second Law of Motion in projections on axis "x" and "y":


"Pcos\theta+F_{fr}-Wsin\theta=0,""N-Wcos\theta-Psin\theta=0."

Let's rewrite our equations:


"Pcos\theta+\mu_sN-Wsin\theta=0""N=Wcos\theta+Psin\theta."

Let's substitute "N" into the previous equation and solve for "P"


"Pcos\theta+\mu_s(Wcos\theta+Psin\theta)-Wsin\theta=0,""P=\dfrac{W(sin\theta-\mu_scos\theta)}{cos\theta+\mu_ssin\theta},""P=\dfrac{2197\ N\cdot(sin45^{\circ}-0.3\cdot cos45^{\circ})}{cos45^{\circ}+0.3\cdot sin45^{\circ}}=1183\ N."